A cruise ship can sail 32 mph in calm water. Sailing with the Gulf Stream, the ship can sail 234 mi in the same amount of time as it takes to sail 150 mi against the Gulf Stream. Find the rate of the Gulf Stream.

Answer by mananth(16949)

A cruise ship can sail 32 mph in calm water. Sailing with the Gulf Stream, the ship can sail 234 mi in the same amount of time as it takes to sail 150 mi against the Gulf Stream. Find the rate of the Gulf Stream.

let the sped of the stream be x

-------------------distance------------ speed------------------ time
with stream---------- 234-----------------32+x-------------------- 23/(32+x)
against stream--------150-----------------32-x----------------- 150/(32-x)
time taken is the same
so 234/(32+x) = 150/(32-x)
234*(932-x) = 150*(32+x)
7488 - 234x = 4800 +150x
7488 -4800= 150x+234x
2688= 384x
x= 7 mph speed of the stream